### v1

# Post-hoc Study of Climate Microtargeting on Social Media Ads with LLMs: Thematic Insights and Fairness Evaluation

Abstract.
Climate change communication on social media increasingly employs microtargeting strategies to effectively reach and influence specific demographic groups. This study presents a post-hoc analysis of microtargeting practices within climate campaigns by leveraging large language models (LLMs) to examine Facebook advertisements. Our analysis focuses on two key aspects: demographic targeting and fairness. We evaluate the ability of LLMs to accurately predict the intended demographic targets, such as gender and age group, achieving an overall accuracy of 88.55%88.$5$. Furthermore, we instruct the LLMs to generate explanations for their classifications, providing transparent reasoning behind each decision. These explanations reveal the specific thematic elements used to engage different demographic segments, highlighting distinct strategies tailored to various audiences. Our findings show that young adults are primarily targeted through messages emphasizing activism and environmental consciousness, while women are engaged through themes related to caregiving roles and social advocacy. In addition to evaluating the effectiveness of LLMs in detecting microtargeted messaging, we conduct a comprehensive fairness analysis to identify potential biases in model predictions. We assess disparities in accuracy and error rates across demographic groups using established fairness metrics such as Demographic Parity, Equal Opportunity, and Predictive Equality. Our findings indicate that while LLMs perform well overall, certain biases exist, particularly in the classification of senior citizens and male audiences. The analysis of thematic explanations uncovers recurring patterns in messaging strategies tailored to various demographic groups, while the fairness analysis underscores the need for more inclusive and unbiased targeting methods. By showcasing the efficacy of LLMs in dissecting and explaining targeted communication strategies and by highlighting fairness concerns, this study provides a valuable framework for future research aimed at enhancing transparency, accountability, and inclusivity in social media-driven climate campaigns.
microtargeting, climate campaigns, social media, demographic targeting, fairness, thematic analysis, large language models
journalyear: 2024copyright: rightsretainedconference: ACM Conference; 2024; N, NN, NNNbooktitle: ACM Conference, 2024, N, NN, NNNccs: Information systems Online advertisingccs: Computing methodologies Natural language processingccs: Applied computing Law, social and behavioral sciences
1. Introduction
Climate change represents one of the most pressing global challenges of the 21st century, necessitating widespread public awareness and engagement to drive meaningful environmental action (Moritz and Agudo, 2013; Dessler and Theater, 1995). As traditional media channels evolve, social media has emerged as a pivotal platform for climate communication, enabling organizations, activists, and policymakers to disseminate information, mobilize support, and influence public discourse on environmental issues (Nosek, 2020; Hestres and Hopke, 2017; Adger et al., 2003). The interactive and targeted nature of social media advertising allows for the customization of messages to resonate with specific audiences, thereby enhancing the effectiveness of communication strategies aimed at fostering climate awareness and behavioral change (Bloomfield and Tillery, 2019; Walter et al., 2018; Stoddart et al., 2016).
In recent years, the utilization of microtargeting strategies in social media campaigns has gained significant traction. Microtargeting involves the precise tailoring of messages to distinct demographic segments based on factors such as age, gender, location, and interests (Prummer, 2020; Hersh, 2015; Barbu, 2014). This approach leverages vast amounts of user data to craft personalized content that is more likely to engage and persuade individual users. In the context of climate communication, microtargeting can enhance the relevance and impact of messages, potentially leading to greater public engagement and support for environmental initiatives. However, the sophistication of these strategies also raises critical questions about the transparency, ethical implications, and overall effectiveness of targeted climate messaging (Islam and Goldwasser, 2024a; Islam et al., 2023b).
Figure 1. Example of climate microtargeting.
Figure 1 illustrates the targeted climate advertisements on social media, with a specific focus on demographic targeting. The first ad source targets young adults aged 18-24, with a message encouraging action against climate change and inviting them to become TCI Ambassadors. The second ad source is tailored to a female audience with a message emphasizing the importance of reducing carbon pollution from home electricity and making a pledge for their children’s future. The third ad source targets a male audience, focusing on the benefits of using clean fuel like Ohio Propane for heating rural homes and fueling appliances.
Despite the growing prevalence of microtargeting in climate campaigns, there remains a limited understanding of the specific techniques and linguistic patterns employed to engage different demographic groups. Traditionally, thematic analysis (TA) has been the preferred method for this task, functioning as a qualitative research technique that focuses on identifying patterns, where themes that emerge from the data drive further analysis (Roberts et al., 2019; Vaismoradi et al., 2016, 2013; Braun and Clarke, 2012, 2006; Tuckett, 2005). However, with the surge of textual data in the digital era, there is an increasing trend towards using computational methods for text analysis (Sivarajah et al., 2017; Jagadish et al., 2014; Fan et al., 2014). TA method often falls short in capturing the nuanced and context-dependent nature of targeted communication. This gap highlights the need for advanced analytical tools that can dissect and interpret the complex language and strategies used in microtargeted climate advertisements. Large language models (LLMs) (Brown et al., 2020), with their robust natural language processing (NLP) capabilities, offer a promising solution to this challenge. In this paper, we investigate whether the newly emerged paradigm in NLP; zero-shot prompting of LLMs (Brown et al., 2020) and the practice of providing explanations of answers — is better equipped to address those challenges.
Explanations are fundamental to human learning (Ahn et al., 1992), as they underscore task principles that facilitate broad generalizations (Lombrozo and Carey, 2006; Lombrozo, 2006). Consider the example text ( “The turbines that provide clean wind energy also create new habitats for fish, keeping the ecosystem healthy.” ) from Figure 2(a) for gender identification task. An explanation can elaborate on a brief answer (e.g., male) by connecting it to the broader reasoning process necessary to solve the problem (e.g., “Emphasizing fish habitats may appeal more to men interested in fishing and environmental conservation.”). Thus, explanations enhance understanding by demonstrating how task principles connect questions to their answers.
While LLMs offer powerful capabilities for analyzing and generating text, their widespread use has also highlighted significant challenges related to fairness (Kumar et al., 2023; Li et al., 2023b, a; Sag, 2023) and bias (Lin et al., 2024; Kotek et al., 2023; Fang et al., 2024; Urman and Makhortykh, 2023; Esiobu et al., 2023). Research has shown that LLMs, like other machine learning (ML) models, can inherit and even amplify biases present in the data they are trained on (Blodgett et al., 2020; Binns, 2018). These biases can manifest in various forms, such as differential accuracy across demographic groups (Dacon and Liu, 2021; Lu et al., 2020; Hitti et al., 2019; D´ıaz et al., 2018), harmful stereotypes (Bianchi et al., 2023; Matthews et al., 2022; Alemany et al., 2022), and discriminatory language patterns (Salguero and Espinilla, 2018). The implications are particularly concerning in high-stakes applications such as healthcare, finance, and social media, where biased outcomes can perpetuate inequities and undermine trust in artificial intelligence (AI) systems (Mehrabi et al., 2021; Buolamwini and Gebru, 2018). Addressing these issues requires fairness evaluations using metrics like Demographic Parity, Equal Opportunity, and Predictive Equality to ensure that models perform equitably across all user groups (Žliobaitė, 2017; Hardt et al., 2016). In the context of this study, examining the fairness of LLMs predictions in demographic targeting is crucial for understanding the microtargeted climate messaging.
In this study, we conduct a post-hoc analysis of climate microtargeting practices on social media by leveraging the power of LLMs (OpenAI’s o1-preview model [^1] https://openai.com/index/learning-to-reason-with-llms/). Post-hoc analysis, typically performed after the main experiment or event, allows us to retrospectively evaluate how effective these campaigns are in targeting specific demographics. Building upon data from previous research by Islam et al. (2023b), we investigate the ability of LLMs to accurately detect targeted messaging based on specific demographic variables, including gender and age group. Additionally, LLMs provide explanations for their classification decisions, offering insights into the thematic and linguistic elements used to engage different audiences. Furthermore, we conduct a comprehensive fairness analysis to identify potential biases in model predictions.
Figure 2. Prompt examples for gender prediction (shown as zero-shot). (a) male, (b) female. Inputs are shown in blue, and outputs are shown in red.
Our exploration leads us to the following research questions (RQ), which are crucial for assessing the potential of LLMs to understand microtargeting patterns and provide deeper and more nuanced insights:
  • RQ1: Given a text, can LLMs predict the targeted demographic of the corresponding text accurately and provide the explanation for the reasoning behind the prediction?
  • RQ2: Which words or phrases are most commonly associated with certain demographics?
  • RQ3: What are the recurrent themes and aspects of explanations provided by LLMs?
  • RQ4: How fair are the LLMs predictions in terms of demographic targeting, and what are the disparities in prediction accuracy and error rates across different demographic groups?
The implications of this research are multifaceted. From an academic perspective, it contributes to the burgeoning field of computational social science (CSS) by showcasing the application of advanced language models in dissecting complex communication strategies. Practically, the findings offer valuable insights for policymakers, environmental organizations, and social media platforms seeking to enhance the transparency, accountability, and inclusivity of their climate communication efforts. By illuminating the specific methods used to tailor messages to different demographics and by highlighting the need for fairer and more inclusive targeting methods, this study lays the groundwork for future investigations into the role of AI in enhancing the efficacy and ethical standards of digital climate advocacy.
2. Related Work
The intersection of microtargeting, social media, and climate communication has drawn considerable interest in recent years, with studies examining how digital platforms influence public opinion and engagement (Islam and Goldwasser, 2024b, a; Holder et al., 2023; Islam et al., 2023b). The growing capabilities of social media platforms to deliver personalized messages based on user data have sparked a significant body of research into computational advertising (Abrams et al., 2022; Huh and Malthouse, 2020; Yang et al., 2017; Zhang et al., 2017), demographic targeting (Allen, 2022; Rummo et al., 2020; Ribeiro et al., 2019; Kaspar et al., 2019; Jansen et al., 2013; Jansen and Solomon, 2010), and the broader implications of these practices for societal discourse.
Microtargeting on social media platforms such as Facebook has been widely studied in the context of political campaigns (Islam et al., 2023a; Capozzi et al., 2021; Serrano et al., 2020; Silva et al., 2020; Capozzi et al., 2020) and public health messaging (Islam and Goldwasser, 2022; Silva and Benevenuto, 2021; Mejova and Kalimeri, 2020), revealing both the potential benefits and ethical concerns associated with this practice. In the realm of climate communication, microtargeting can be a powerful tool for enhancing message relevance and impact by tailoring content to the specific values and interests of diverse demographic groups (Islam et al., 2023b). However, the effectiveness and ethical implications of such targeted messaging remain under-explored, particularly in terms of how different demographic groups are engaged and whether biases are present in the targeting strategies.
Recent advances in LLMs have demonstrated their capability for in-context learning (ICL), significantly enhancing their ability to perform tasks traditionally handled by humans (Chowdhery et al., 2023; Kojima et al., 2022; Le Scao et al., 2022; Brown et al., 2020; Gilardi et al., 2023; De Paoli, 2023; Dai et al., 2023; Chiang and Lee, 2023; Ziems et al., 2024). This progress suggests a strong potential for effectively applying LLMs to our specific task. In the realms of qualitative research (QR) and NLP, innovative methods are being explored to integrate LLMs into TA. Researchers have proposed various frameworks, including an LLM-in-the-loop approach (Islam and Goldwasser, 2024b, a; Dai et al., 2023), integrating GPT-3 with expert-designed codebooks (Xiao et al., 2023), and developing collaborative interfaces that utilize LLMs for code generation and support in decision-making processes (Gao et al., 2023). Other recent work has shown that LLMs can benefit from examples that decompose the reasoning process (can be seen as an explanation), leading to an answer (Wei et al., 2022). Despite the impressive capabilities of LLMs, there are concerns about fairness, accountability, and transparency in LLMs predictions which have been highlighted in recent literatures (Anthis et al., 2024; Dai et al., 2024; Kumar et al., 2023; Li et al., 2023b; Wu et al., 2023; Bender et al., 2021), emphasizing the need for rigorous evaluation of biases and disparities in model performance across different groups.
In this paper, we leverage zero-shot capabilities of LLMs to identify targeted demographics and provide explanations of demographic targeting regarding climate-related advertisements on Facebook. Besides, we develop a new set of themes and aspects based on those explanations specifically tailored for analyzing messaging. Additionally, we extend previous research by not only focusing on the accuracy of these predictions but also conducting a comprehensive fairness analysis to identify and address potential biases in the model’s performance.
3. Dataset
We investigate the climate campaigns case study for this work. We work on the corpus of 213722137221372 ads released by Islam et al. (2023b). This dataset includes climate-related English ads on Facebook from the US, spanning from January 202120212021 to January 202220222022. Each ad includes the following attributes: ad ID, title, ad description, ad body, funding entity, spend, impressions, and distribution of impressions broken down by gender (male, female, unknown), age (seven groups), and location down to the state level in the USA. Additional details about the dataset can be found in the original publications.
For this work, we consider two demographic indicators, i.e., gender and age group. We consider two gender categories, i.e., male vs female. Regarding age group, we consider four age group categories, i.e., young adults whose age range is 18-24, early working age group (25-44), late working age group (45-64), and senior citizens (65+). We found approximately 227227227 targeted unique ads; among them, 595959 ads targeting only females and 474747 ads targeting only males. However, we find 252525 ads target only young adults, 828282 ads target only the early working age group, 888 ads target only the late working age group, and 666 ads target only senior citizens.
Figure 3. Prompt template for targeted demographic prediction (shown as zero-shot). (a) gender, (b) age group. Inputs are shown in blue, and outputs are shown in red.
4. Task Definition
The identification of the targeted demographic (with explanation) in a text involves the following steps:
Gender prediction with Explanation: For a given text $t$, the task involves identifying the targeted gender and explaining the rationale behind its selection.
Age group prediction with Explanation: Subsequently, the task requires predicting the targeted age group and providing an explanation for the specific choice.
To predict the targeted demographics, we employ zero-shot prompting using the most recently [^2] September 12, 2024 released by OpenAI, o1-preview model [^3] https://openai.com/index/introducing-openai-o1-preview/. This is a new large language model that uses reinforcement learning (RL) and chain of thought (COT) techniques for complex reasoning, allowing it to think through a detailed internal process before responding to users.
The prompt template for the demographic prediction task with an explanation using LLMs can be found in Figure 3. Figure 2 shows the example prompts for gender prediction. Figure 5 (in App. A) shows the example prompts for age group prediction from the climate campaign dataset.
5. Experimental Setup
In this work, we use OpenAI Playground API to run o1-preview by keeping the default parameters.
Category Total Ads Correct Pred. Acc. (%) Misclass.
All 227 201 88.55 -
Female 59 56 94.92 3 (Male)
Male 47 40 85.10 7 (Female)
Young adults 25 22 88.00 2 (Early Working), 1 (Late Working)
Early Working 82 75 91.46 4 (Young), 4 (Late Working)
Late Working 8 6 75.00 2 (Early Working)
Senior 6 2 33.33 3 (Young), 1 (Late Working)
Table 1. Accuracy and Misclassifications for Demographics.
5.1. Results
[…]
For baseline comparison, we use open sourced LLMs Llama 3 (llama3-70b-8192444https://github.com/meta-llama/llama3) (Touvron et al., 2023) and Mistral Large 2 (mistral-large-2407555https://mistral.ai/news/mistral-large-2407/) (Jiang et al., 2023). OpenAI’s o1-preview model outperforms the baselines both in gender and age group predictions (Table 2).
Model Demo. Acc. (%) Macro Avg. F1 (%)
o1-preview gender 90.57 90.35
o1-preview age 85.95 71.00
llama 3 gender 80.19 79.67
llama 3 age 58.68 36.84
Mistral Large 2 gender 82.08 82.07
Mistral Large 2 age 74.38 48.68
Table 2. Baseline comparisons.
[…]
Early Working: High precision (0.950.950.95) and recall (0.900.900.90), indicating strong performance in identifying individuals in this age group.
Late Working: Moderate precision (0.500.500.50) and high recall (0.750.750.75), suggesting that while the model captures most of the actual instances, it has a higher rate of false positives.
Senior: Perfect precision (1.001.001.00) but low recall (0.330.330.33), meaning the model is very accurate when it predicts the Senior class but misses a significant number of actual senior instances.
Young: Good precision (0.760.760.76) and high recall (0.880.880.88), showing reliable performance in identifying younger individuals.
5.2. Error Analysis
Table 10 in the App. B.3 presents an analysis of ad misclassifications based on gender and age group predictions. Each entry includes the actual demographic, the predicted demographic, and a brief explanation generated by LLMs. Table 10 and explanations highlight how specific patterns and themes within an ad can lead to demographic misclassifications. In some cases, traditional gender roles and age-related interests played a significant role in these misclassifications. Understanding these nuances can help in refining predictive models and improving the accuracy of demographic targeting in future ad campaigns.
Figure 4. Prompt template for generating theme and aspects from predictions and explanations (shown as zero-shot). Inputs are shown in blue, and outputs are shown in red.
6. Ad Content Analysis
To determine which words or phrases are most commonly associated with certain demographics (answering RQ2), we identify the top-5 most frequent bigrams (two-word pairs) and trigrams (three-word pairs) of the ad content for male, female, young adults, and early working age group. Due to the small sample size, we do not show this analysis for the late working age group and senior citizens. Table 11 in App. C details the results.
These frequent bigrams and trigrams highlight key themes in the ads, particularly focusing on the phrases related to climate action, urgency (emergency), and leadership for young adults. On the other hand, frequent bigrams and trigrams of early working age group highlight the central themes of climate change, clean energy, and calls to action in the ads. Moreover, for male, these frequent bigrams and trigrams highlight the recurring themes related to climate change, clean energy, and specific campaigns or initiatives in the ads. In contrast, for female, these frequent bigrams and trigrams emphasize recurring themes around climate change, clean energy, and the ‘Build Back Better’ agenda in the ads (Table 11 in App. C ).
To calculate the statistical significance, we perform Chi-Square Test (Cochran, 1952) for Independence between male and female demographics and their bigrams/trigrams. The p-values for both the bigrams and trigrams are greater than the common significance level of 0.050.050.05. This suggests that there is no statistically significant association between gender (male and female) and the frequency of the bigrams and trigrams analyzed. Therefore, we fail to reject the null hypothesis and conclude that the observed differences in bigram and trigram frequencies between males and females could be due to chance. The same trend is noticed while performing a Chi-Square Test for Independence between the young adults (18-24) and early working (25-44) age groups and their bigrams/trigram.
Gender Theme of Explanation Aspects of Explanation
Male Perceived Interests and Roles Interest in Technology and Innovation Focus on Economic and Financial Issues Property and Land Management Traditional Male Activities Engagement in Political and Infrastructure Topics Conservative Views and Skepticism
Female Roles as Caregivers, Environmental Advocates, and Socially Conscious Individuals Parental and Caregiving Roles Environmental Consciousness Social Welfare and Community Involvement Empathy and Emotional Appeal Female Empowerment and Leadership Health and Safety Concerns
Table 3. Gender based Themes and Aspects of Explanations.
7. Thematic Insights of Explanations
As LLMs provide explanations to provide reasoning behind their prediction, we use those explanations for thematic analysis to answer RQ3. In this analysis, we only include the correct predictions and their explanations.
7.1. Themes and Aspects of Gender Explanations
We prompt LLMs in a zero-shot manner to provide the common theme and aspects under specific theme of the explanations from 404040 correct male predictions and 565656 correct female predictions. The prompt template is shown in Figure 4. We detail the theme of the gender explanation and aspects of the explanation in Table 3.
From $1^{st}$ row of Table 3, we notice that the overall theme revolves around targeting men by aligning ads with their perceived interests and roles in technology, finance, property, traditional male activities, and political or economic discourse. The explanations consistently emphasize the following aspects:
Interest in Technology and Innovation: Men are often depicted as being more engaged with technology, engineering, and renewable energy solutions. Ads related to technical aspects of engines, energy efficiency, and infrastructure are considered more likely to appeal to men.
Focus on Economic and Financial Issues: Many explanations suggest that men are more concerned with economic benefits, investment opportunities, and financial savings, making them the likely target for ads that emphasize these aspects.
Property and Land Management: The theme of land ownership, property value improvement, and land management is frequently mentioned, with the assumption that men are more interested in these areas.
Traditional Male Activities: Ads that involve traditionally male-oriented activities, such as beer consumption, vehicle-related savings, home maintenance, and physical strength, are seen as more likely to target men.
Engagement in Political and Infrastructure Topics: Men are often portrayed as more engaged in political discourse, infrastructure initiatives, and discussions around energy and policy, making them the primary audience for ads focused on these themes.
Conservative Views and Skepticism: Some explanations suggest that men are more likely to resonate with conservative views, skepticism about environmental claims, and anti-establishment sentiments.
From $2^{nd}$ row of Table 3, we observe that the overall theme revolves around targeting women by aligning ads with their roles as caregivers, environmental advocates, and socially conscious individuals who prioritize the well-being of their families, communities, and the environment. The explanations consistently emphasize the following aspects:
Parental and Caregiving Roles: Many explanations highlight that women, particularly mothers, are more likely to resonate with messages about protecting children’s futures, parental responsibilities, and family well-being. These ads often appeal to maternal instincts and the role of women as primary caregivers.
Environmental Consciousness: Women are frequently depicted as being more engaged with environmental issues, sustainability, and community health. The explanations suggest that women are more proactive and vocal about climate change, conservation, and eco-friendly initiatives.
Social Welfare and Community Involvement: The explanations note that women are more likely to be concerned with social issues such as paid leave, affordable childcare, healthcare, and community well-being. Ads that emphasize these themes are seen as more likely to appeal to women.
Empathy and Emotional Appeal: The explanations often mention that women are more responsive to ads that evoke empathy, emotional concerns, and collective action. This includes ads that focus on protecting the environment for future generations and supporting social safety nets.
Female Empowerment and Leadership: Some explanations specifically mention themes of women’s empowerment, leadership, and support for female scientists or leaders. These themes are likely to resonate more with female audiences who identify with or support gender equality and empowerment.
Health and Safety Concerns: Women are portrayed as being more attentive to issues related to health, safety, and the well-being of their families and communities. This includes a strong focus on environmental health and sustainability.
7.2. Themes and Aspects of Age Explanations
We prompt LLMs in a zero-shot manner to provide the common theme and aspects under specific theme of the explanations from 222222 correct young adult predictions, 757575 correct early working age group predictions, 666 correct late working age group predictions, and 222 correct senior citizen predictions. We detail the theme of the age group explanation and aspects of that explanation in Table 12 in App. D.
From $1^{st}$ row of Table 12 (App. D), we observe that the overall theme revolves around the activism and the environmental consciousness of young adults, positioning them as a key demographic for campaigns and initiatives focused on climate change and sustainability. The explanations consistently highlight the following aspects:
Passion for Climate Action: Young adults are described as being particularly passionate about addressing climate change, often leading or participating in environmental activism and campaigns.
Support for Bold Environmental Leadership: This age group is likely to support bold and urgent actions related to environmental protection and sustainability.
Engagement with Activism: The explanations emphasize that young adults are more likely to be engaged in climate-related activism and are motivated to take meaningful actions.
Desire for Immediate Change: There is a recurring mention of the desire for immediate and meaningful change, reflecting the urgency with which young adults approach environmental issues.
Participation in Training and Advocacy: The group is also characterized as eager to participate in training programs and initiatives that allow them to contribute actively to environmental causes.
From $2^{nd}$ row of Table 12 in App. D, we can see that the overall theme revolves around the proactive and responsible mindset of early working-age adults, who are not only financially capable but also motivated by a strong sense of social and environmental responsibility. They are seen as key targets for initiatives that combine sustainability with practical, career-oriented, and family-focused benefits. The explanations consistently emphasize the following aspects:
Environmental Consciousness: This age group is described as being highly engaged with environmental issues, such as climate change, sustainability, and clean energy. They are likely to support initiatives and products that align with eco-friendly values.
Financial Stability and Disposable Income: Many explanations note that individuals in this group have disposable income, making them financially capable of supporting and investing in sustainable products, services, and causes.
Parental and Future Concerns: This demographic is often portrayed as parents or future-focused individuals who are concerned about the impact of environmental issues on their children and future generations.
Career Engagement and Professional Roles: The explanations frequently mention that this age group is active in their careers, often holding decision-making roles that influence corporate and household sustainability practices.
Interest in Innovation and Technology: Individuals in this age group are also depicted as being interested in innovative industries, clean energy solutions, and sustainability technologies, which align with their professional and personal goals.
Social and Political Engagement: The group is characterized as being engaged in socio-political issues, particularly those related to corporate accountability, sustainability, and environmental advocacy.
From $3^{rd}$ row of Table 12 in App. D, we can notice that the overall theme revolves around the responsibilities and concerns of individuals in the late working (45-64) age group, focusing on their roles as homeowners, voters, and economically engaged citizens who are likely to be influenced by environmental, economic, and policy-related messaging. The explanations specifically emphasize the following aspects:
Economic and Environmental Responsibility: Many of the explanations mention that individuals in this age group are concerned with sustainability, home energy efficiency, and environmental impact. They are likely to invest in public resources and adopt changes that contribute to economic and environmental sustainability.
Homeownership and Financial Stability: This demographic is characterized as established homeowners who are financially secure. They are seen as key targets for changes related to home energy efficiency, such as adopting solar power, due to their financial means and homeownership status.
Voter and Policy Engagement: The explanations suggest that this age group is politically active, particularly concerned with public safety, and likely to support policy changes by voting on local measures.
Economic Concerns: There is an emphasis on economic factors such as unemployment, inflation, and gas prices, with concerns about current economic policies affecting their businesses and financial stability.
From $4^{th}$ row of Table 12 (App. D), we notice that the overall theme centers on health and safety concerns that are particularly important to senior citizens, with a focus on programs that cater to their specific needs and the heightened risks they face in certain situations. The key aspects highlighted in the explanations are:
Health and Wellness Programs: The first explanation mentions programs like SilverSneakers and Silver&Fit, which are specifically designed for senior citizens to support their physical health and well-being.
Vulnerability and Safety: The second explanation focuses on the increased vulnerability of seniors to COVID-19 and emphasizes the risks they face, particularly in the context of political decisions or public health issues.
8. Fairness and Bias Analysis
In this section, we present a comprehensive fairness analysis of the model for gender and age group classifications to answer RQ4. We evaluate the models using established fairness metrics such as Demographic Parity, Equal Opportunity, and Predictive Equality to assess their performance across different groups. By analyzing confusion matrices and classification reports, we identify any disparities in prediction accuracy and error rates between groups. Our analysis aims to identify biases, investigate the underlying reasons for any observed biases. The insights gained from this analysis are critical for guiding future research in developing fair and inclusive algorithms.
Predicted Gender Demographic Parity Ratio
Female 1.0678
Male 0.9149
Table 4. Demographic Parity for Gender.
Predicted Age Group Demographic Parity Ratio
Early Working 0.95
Late Working 1.50
Senior 0.33
Young 1.16
Table 5. Demographic Parity for Age Group.
Age Group Equal Opportunity (TPR) Predictive Equality (FPR)
Early Working 0.90 0.10
Late Working 0.75 0.05
Senior 0.33 0.00
Young 0.88 0.07
Table 6. Equal Opportunity (True Positive Rate) and Predictive Equality (False Positive Rate) for Age Groups.
Ad Text Prediction Explanations by LLMs
The Hebrew University Center for Climate Science was established in Israel to fight climate change worldwide. Young Targets university-age students interested in climate science to study and combat climate change.
Tell Rep. Schrader: Now is the time to go big on climate. VOTE YES on the Build Back Better Act. Young Targets young adults passionate about climate action and eager to influence political decisions for their future.
Get clean water and pollution-free electricity to all of America. Young Targets environmentally conscious young adults concerned about sustainability and future impact of clean water and energy.
Let’s make one thing clear: Pennsylvania will be the single most competitive Senate race of 2022… It’s one of the reasons I am running for the U.S. Senate seat in Pennsylvania. Late working Targets 45-64-year-olds by focusing on working families, anti-celebrity politics, and referencing the Trump Era.
Table 7. Misclassified Senior Instances.
8.1. Fairness Analysis on Gender Prediction
To assess the fairness of the model on gender prediction, we compute several fairness metrics, including Demographic Parity, Equal Opportunity, and Predictive Equality.
Demographic parity examines whether each gender group receives positive predictions at equal rates. Table 4 presents the Demographic Parity ratios for each gender. A ratio of 111 indicates perfect parity. The results show that females have a slightly higher likelihood of receiving positive predictions compared to males, suggesting a minor imbalance favoring the female class.
Equal Opportunity (True Positive Rate) focuses on the True Positive Rates (TPR) across gender groups, measuring the model’s ability to correctly identify positive instances within each group. We have Female TPR: $\mathbf{0.95}$ and Male TPR: $\mathbf{0.85}$. The TPR for females is higher by 0.100.100.10, indicating that the model is more effective at correctly identifying females than males.
Predictive Equality (False Positive Rate) assesses the False Positive Rates (FPR) across gender groups, reflecting the rate at which negative instances are incorrectly labeled as positive. We achieve Female FPR: $\mathbf{0.07}$ and Male FPR: $\mathbf{0.05}$. The slightly higher FPR for females suggests that females are more likely to be incorrectly predicted as positive compared to males.
8.2. Fairness Analysis on Age Group Prediction
Results in Table 5 show that the late working group has a higher likelihood of receiving positive predictions, while the senior group has a significantly lower rate, indicating potential bias against seniors. The senior group has a Demographic Parity ratio of 0.330.330.33, significantly lower than the ideal value of 111, suggesting under representation in positive predictions.
Table 6 shows that the TPR for the senior group is notably lower, suggesting the model is less effective at correctly identifying individuals in this age group. The senior group has an FPR of 0, indicating no false positives, while the early working group has the highest FPR among the groups.
To understand the observed bias in the misclassified senior age group, we conduct an analysis of the misclassified instances (Table 7). The goal is to identify patterns and underlying reasons why the model predicts seniors as belonging to the young and late working age group. The misclassification of the Senior age group as Young or Late Working can be attributed to several factors:
Thematic Content and Topic Association: The first three misclassified ads (Table 7) focus on climate change, environmental activism, and sustainability. These topics are often associated with younger demographics, particularly young adults (18-24), who are perceived to be more engaged in activism and environmental causes. The model appears to have learned an association between these topics and the Young age group, leading to misclassification when seniors engage with similar content.
In Table 7, the fourth instance shows a misclassification where the model predicts it as a late working (45-64 years) age group. The ad mentions ‘working families’, which is a term commonly associated with individuals in the Late Working age group who are actively engaged in the workforce and supporting families. The content revolves around a political campaign emphasizing the need for change and active participation, themes often associated with the 456445-6445 - 64 age demographic who are typically more politically active and influential.
Lack of Age-Specific Cues: The misclassified ads do not contain explicit references to seniors or age-specific language that would signal the content is intended for the senior demographic. The language is broad and does not mention age-related concerns, such as retirement, health issues prevalent among seniors, or senior-specific programs.
Reliance on Stereotypical Associations: The explanations generated by LLMs indicate that the model relies on stereotypes, associating certain topics exclusively with specific age groups. By assuming that environmental activism is primarily of interest to young adults, the model overgeneralizes by overlooking the possibility that seniors are also engaged in these issues.
Feature Representation Limitations: The model may lack features that capture subtle cues indicating the ad’s target age group when explicit age markers are absent. The model may not effectively utilize contextual information that could hint at the intended audience beyond topic associations. For example, the interests and concerns of the late working and senior age groups can overlap, especially in areas like politics and social change.
9. Conclusion
Our work advances the discourse on climate microtargeting by demonstrating the utility of LLMs in accurately detecting and explaining targeted messaging strategies on social media. We provide a broader theme and various aspects under each theme based on the explanations from LLMs, which we hope will be an important contribution to the CSS community. Besides, the fairness analysis conducted in our study underscores the importance of evaluating and addressing the biases. Disparities in prediction accuracy and error rates, particularly in underrepresented groups, highlight the need for more inclusive and equitable targeting methods. Ultimately, this study lays the groundwork for future investigations into the role of AI in enhancing the efficacy, transparency, and accountability of digital climate advocacy.
10. Limitations
Our analysis relies on OpenAI o1-preview model. We chose o1-preview instead of the open-source counterparts due to computational resource constraints. We show our analysis on climate campaigns dataset, but our approach can easily be adapted in any dataset. It is designed to be scalable without any modifications.
11. Ethics Statement
To the best of our knowledge, we did not violate any ethical code while conducting the research work described in this paper. We report the technical details for the reproducibility of results. In this paper, we did not introduce any new dataset; instead, we experimented using existing dataset that are adequately cited. The author’s personal views are not represented in any qualitative results we report, as it is solely an outcome derived from a machine learning model.
Appendix A Prompting
Figure 5 shows the example prompts for age group prediction from the climate campaign dataset.
Figure 5. Prompt examples for age group prediction (shown as zero-shot). (a) young adults (18-24), (b) early working age (25-44), (c) late working age (45-64). Inputs are shown in blue, and outputs are shown in red.
Appendix B Experiment Details
In this section, we provide experimental results.
B.1. Confusion Matrix
Figure 6 shows the confusion matrices for the targeted demographic group. Figure 6(a) shows the confusion matrix for gender, and Figure 6(b) shows the confusion matrix for age group prediction.
(a) gender.
(b) age
Figure 6. Confusion matrices.
B.2. Demographic Classification Report
Table 8 provides the classification report with detailed performance metrics for each gender class. Table 9 shows the classification report with detailed performance metrics for each age group.
Gender Precision Recall F1-Score Support
Female 0.89 0.95 0.92 59
Male 0.93 0.85 0.89 47
Accuracy 0.91 106
Macro Avg 0.91 0.90 0.90 106
Weighted Avg 0.91 0.91 0.91 106
Table 8. Classification Report for Gender.
Age Group Precision Recall F1-Score Support
Early Working 0.95 0.90 0.92 82
Late Working 0.50 0.75 0.60 8
Senior 1.00 0.33 0.50 6
Young 0.76 0.88 0.81 25
Accuracy 0.86 121
Macro Avg 0.80 0.72 0.71 121
Weighted Avg 0.88 0.86 0.86 121
Table 9. Classification Report for Age Group.
Ad Text Actual Prediction (Misclassification) Explanations by LLMs
Meet the first carbon negative fragrance. Air Eau de Parfum converts CO2 into something beautiful: sustainable ethanol blended with notes of orange peel, fig leaf, powdery musk and tobacco invigorating scent that redefines what is possible. Available exclusively at aircompany.com Male Female The fragrance notes like fig leaf, powdery musk, and sustainable ethanol typically resonate more with female audiences in fragrance marketing.
Over generations, Illinois farm families have learned how to work with our climate to produce healthy crops and livestock. Female Male The ad emphasizes farming and agriculture, traditionally male-dominated fields, making men the more likely target audience for this message.
Renewable energy is projected to surpass fossil gas as the dominant source of energy production by 2030. Texas, a leader in wind output, is in a great position as the nation transitions. Young Early working This group is likely investing in sustainable energy options, driven by long-term economic and environmental considerations.
Take Action Now to Fight Climate Change! Plant a tree in California treestoglobe.org/plantingchallengecalifornia Early working Young This age group is often more environmentally active and responds well to social media campaigns promoting climate action initiatives.
The worst impacts of climate change could be irreversible by 2030. The time to switch to a renewable energy plan is now. Not tomorrow. Not next week. Late working Early working They can switch energy plans and are motivated to act before irreversible impacts by 2030.
Tell Rep. Schrader: Now is the time to go big on climate. VOTE YES on the Build Back Better Act. Senior Young Targets young adults passionate about climate action and eager to influence political decisions for their future.
Table 10. Error Analysis.
B.3. Demographic Misclassifications
Table 10 presents an analysis of ad misclassifications based on gender and age group predictions.
Appendix C Bigrams and Trigrams
Table 11 shows the top-5 most frequent bigrams (two-word pairs) and trigrams (three-word pairs) of the ad content for male, female, young adults, and early working age groups.
Demo. Indicator Groups Top 5 Bigrams: # Occurrences Top 5 Trigrams: # Occurrences
Young adults (18-24) “bold climate” - 10 occurrences “climate emergency” - 10 occurrences “declare climate” - 10 occurrences “climate leaders” - 6 occurrences “emergency need” - 6 occurrences “declare climate emergency” - 10 occurrences “bold climate leaders” - 6 occurrences “climate emergency need” - 6 occurrences “climate leaders protect” - 6 occurrences “emergency need bold” - 6 occurrences
Age group Early working (25-44) “climate change” - 10 occurrences “fight climate” - 7 occurrences “clean energy” - 6 occurrences “america clean” - 3 occurrences “energy future” - 3 occurrences “fight climate change” - 7 occurrences “america clean energy” - 3 occurrences “clean energy future” - 3 occurrences “california treestoglobe org” - 2 occurrences “fueling america clean” - 2 occurrences
Male “climate change” - 5 occurrences “clean energy” - 4 occurrences “10 million” - 3 occurrences “million trees” - 3 occurrences “carbon emissions” - 2 occurrences “10 million trees” - 2 occurrences “carbon emissions 50” - 2 occurrences “clean energy corridor” - 2 occurrences “don live way” - 2 occurrences “torched earth ale” - 2 occurrences
Gender Female “climate change” - 7 occurrences “build back” - 6 occurrences “back better” - 6 occurrences “clean energy” - 5 occurrences “protect your” - 4 occurrences “build back better” - 6 occurrences “protect your kids” - 4 occurrences “affordable child care” - 3 occurrences “fighting climate change” - 3 occurrences “climate change and” - 3 occurrences
Table 11. Top-5 Most Frequent Bigrams and Trigrams Across Different Demographic Indicators and Groups.
Appendix D Thematic Analysis of Age Group Explanation
Table 12 details the theme of the age group explanation and aspects of that explanation.
Age group Theme of Explanation Aspects of Explanation
Young adults (18-24) Activism and Environmental Consciousness Passion for Climate Action Support for Bold Environmental Leadership Engagement with Activism Desire for Immediate Change Participation in Training and Advocacy
Early working (25-44) Proactive and Responsible Mindset Environmental Consciousness Financial Stability and Disposable Income Parental and Future Concerns Career Engagement and Professional Roles Interest in Innovation and Technology Social and Political Engagement
Late working (45-64) Responsibilities and Concerns Economic and Environmental Responsibility Homeownership and Financial Stability Voter and Policy Engagement Economic Concerns
Senior (65+) Health and Safety Concerns Health and Wellness Programs Vulnerability and Safety
Table 12. Age group based Themes and Aspects of Explanations.